# Adapting OpenAI’s CLIP Model for Few-Shot Image Inspection in Manufacturing Quality Control: An Expository Case Study with Multiple Application Examples

### v1

@@ 7.1 Dataset Description
Our fourth case study examines the recently published Renault pipe staple dataset (6), which presents a real-world automotive assembly inspection challenge. The dataset consists of images capturing the assembly of flexible cables secured by metallic rectangular clamps in automotive manufacturing. Figure 9 shows representative examples of nominal and defective cases from the dataset.
@@ 8.1 Dataset Description
Manual evaluation by human experts introduces significant limitations in industrial practice (15). The growing volume of material characterization data makes manual labeling resource-intensive. Expert analysis inherently contains subjective biases, limiting the detection of subtle variations within natural data variance. Therefore, the materials science and manufacturing communities have increasingly adopted automated classification methods for microstructure analysis (9, 26). Our investigation of this case is motivated by the suitability of our proposed few-shot learning approach with CLIP to applications with limited training data. If successful, our proposed approach can automate the inspection of microstructure images to reduce the cost associated with expert-driven microstructural measurements.
We simulated EBSD microstructure images by combining random Voronoi tessellation with crystal orientation distribution functions (ODFs) (1). Voronoi edges represent grain size and morphology. The ODF assigns crystal orientations to grains based on predefined probability functions. Crystal orientations map to colors using standard EBSD measurement schemes. Figure 10 illustrates the four example simulated microstructures investigated in our study: (a) uniform (nominal) with random crystal orientations from a uniform distribution, (b) clustered orientation showing a band of grains with predominant crystal orientation surrounded by uniform ODF grains, (c) bimodal with crystal orientations from “bimodal” distribution, and (d) single crystal-like where crystals align toward one orientation. These classes reflect EBSD patterns in metal additive manufacturing. The uniform ODF represents nominal conditions. The other defective classes indicate defects from local or global heating and cooling dynamics variations during manufacturing.
@@
Our ODF simulations used weighted schemes to generate three severity levels per class (“low”, “medium”, “high”). These levels represent different degrees of deviation from the nominal uniform ODF. The simulated microstructural images are 400 x 400 pixels. For technical implementation details, we refer the interested reader to (31).
@@ 8.2 Few-Shot Classification Performance
Similar to case studies 1-4, Table 9 presents the performance of the few-shot learning for the binary image classification experiment (nominal vs. defective). Here, we define an image as defective if it belongs to any of the three primary defective classes (“band”, “bimodal”, and “single-crystal”). This few-shot learning binary classification results show that the model excelled at detecting defects with a sensitivity of 0.987, i.e., 222 out of the 225 defective images were correctly classified. Furthermore, nominal images were correctly classified as nominal 80.9% of the time, with 182 out of 225 nominal images classified correctly. Due to the balanced nature of the dataset, the overall accuracy represents the average of those two numbers and is approximately 90%.
@@
Given the relatively large number of potential defect classes (up to nine if we consider the class–severity combination), it would be interesting to investigate the model’s multi-class classification performance. Note that we can extract its multiclass classification performance without the need to “retrain” or “rerun” the model. Recall that our approach takes every test image and computes its cosine similarity with every image within the “nominal” and “defective” training dataset. From this comparison, we identify the most relevant image within each nominal and defective class, their match probability, and training-based captions (detailed image description) for its most similar training “nominal” and “defective” images. Consider our test set image microstructure_039.png to explain this concept further. The output from our CLIP-based implementation is shown in Figure 11(a). The Figure shows that the overall classification result is defective since the probability of being defective (0.505) is slightly larger than non-defective (0.495). We can also see that its most similar defective image is microstructure_001.png, which had a description/caption that ends with “low single crystal structure”. We extracted this description using simple text extraction strategies to obtain our multi-class results. We utilized two extraction strategies: (a) focusing only on whether it is “band”, “bimodal”, or “single-crystal”, and (b) assigning it to one of the nine defective sub-labels shown in Figure 10 by extracting severity (“low”, “medium”, and “high”) and the primary defect class (“band”, “bimodal”, and “single-crystal”) indicators.
@@
Figure 12: Confusion matrices and the corresponding classification performance metrics for primary and secondary multi-class labels for Case Study 5.
@@
The second row in Figure 12 captures the confusion matrix for the nominal class and the nine defective sublabels alongside its corresponding classification metrics table. As expected, with the increase in the classification degree of difficulty (the model must get both the primary defect class and its severity correct), the overall accuracy drops to 74.4%. However, compared to the primary classification example, the decrease in performance can only be attributed to misclassification errors among different severities of the same type of defect (crystal orientation) rather than errors between one crystal orientation and another.
@@
These results highlight the proposed approach’s capability to classify complex microstructure data through few-shot classification. The results also show the simplicity of extending our binary classification approach to multiclass classification without rerunning the model. This allows practitioners to perform classifications sequentially: first distinguishing between nominal and defective data and then identifying specific types of defects. Note that our results for this case study will likely be improved if we increase the number of learning examples, as shown in case studies 2 and 3. However, we do not explore this here for three reasons: (a) the obtained results demonstrate the approach’s potential and were satisfactory; (b) our primary rationale for presenting this case study was to show the ease of translating the binary classification results into a multi-class classification output, without the need to rerun or “retrain” the model; and (c) conciseness.
@@ 9.1 Insights from the Experiments
Our case studies demonstrate that CLIP, when adapted for few-shot learning, can serve as a powerful yet simple baseline for image-based quality control. The model consistently performed well with relatively small learning sets across most applications. For instance, in the metallic pan case study, we achieved 94% accuracy with just 10 examples per class, while the stochastic textured surfaces case demonstrated 97% accuracy with 50 examples per class. Even in the more challenging extrusion profile case, the model reached 80% accuracy with 350 examples per class, despite working with very low-resolution ($85\times 85$ pixel) images. This dataset is still much smaller than the $10,000+$ examples used in deep learning applications (e.g., see 13).
CLIP successfully processed images ranging from $250\times 250$ pixels (STS case) to full-resolution $3264\times 2448$ images (metallic pan case). While the ViT-L/14 model is designed for $336\times 336$ pixel inputs, it can effectively handle smaller and larger images through appropriate preprocessing. The impact of resolution on performance appears application-dependent, with optimal results generally achieved when image dimensions are larger or relatively close to the model’s design specifications. The results were acceptable in the extrusion case ($85\times 85$ pixels) with 350 few-shot examples per case.
Performance varied significantly based on the complexity of the image. The model excelled with single-component images, as demonstrated in the metallic pan. Particularly noteworthy was its success with stochastic textured surfaces, a traditionally challenging domain for automated inspection. This suggests that transformer-based architectures may inherently capture the statistical properties that define such surfaces, offering a potential new paradigm for texture analysis in manufacturing. Performance in the extrusion case was acceptable with 350 nominal examples. However, the model struggled with multi-component scenarios, as evidenced by the pipe staple case, indicating a limitation in handling complex scenes with multiple interacting elements.
@@
The relationship between learning set size and performance was nonlinear across all cases. For some performance metrics, initial improvements were often large, with substantial gains typically observed in the first 50-100 per class examples. However, diminishing returns appeared after 150-200 per class examples, though the specific inflection point varied by application. This pattern suggests that practitioners can often achieve satisfactory performance with relatively modest data collection efforts, especially when working with single-component or texture-based applications. Not surprisingly, the performance gains across learning set size were nonmonotonic. The variation across learning set size was not likely due to the model’s classification variation; i.e., the same image was classified consistently for a given learning set size. The dips and peaks in performance are likely due to the addition of learning set images, with some additional learning cases making it easier (or harder) to classify the test images.
@@ 10 Concluding Remarks
Our expository study demonstrates that our CLIP-based few-shot learning approach can serve as an effective baseline for image-based quality control in manufacturing. Through five case studies, we showed that CLIP can achieve high classification accuracy with relatively small learning sets, often requiring only 50-100 examples per class compared to the thousands typically needed for deep learning approaches. The microstructure case study further highlighted the model’s ability to adapt to multi-class classification tasks without retraining, demonstrating promising performance across multiple levels of classification granularity (from nominal vs. defective images to identifying specific defect types and their severities). Overall, the CLIP-based few-shot learning approach excelled at single-component inspection, texture analysis, and microstructure classification. However, its performance was poor in complex multi-component scenes, where more advanced computer vision and/or statistical monitoring approaches were needed. Our findings establish CLIP-based few-shot learning as a practical first baseline for quality engineers and researchers. It offers a simple yet powerful solution that should be evaluated before pursuing more complex implementations.
@@ Acknowledgement
The work of Hongyue Sun is partially supported by the NSF Grant No. FM-2134409 and CMMI-2412678.

### v2

@@ 7.1 Dataset Description
Our fourth case study examines the recently published Renault pipe staple dataset (7), which presents a real-world challenge in automotive assembly inspection. The dataset consists of images capturing the assembly of flexible cables secured by metallic rectangular clamps in the automotive manufacturing process. Figure 9 shows representative examples of nominal and defective cases from the dataset.
@@ 7.3 Few-Shot Classification Performance
7.4 Comparison with Other Benchmark Approaches
Our CLIP- and MobileNetV2 few-shot models each used all 100 images for learning. AutoML was trained on 90 images with 10 held out for validation. Table 12 shows that AutoML attains perfect performance across every metric. This reflects its strength when classes are easily separable visually (13). The MobileNetV2-FS prototype achieves 0.750 accuracy, 0.720 sensitivity, 0.780 specificity, and an F1-score of 0.742 in just 1.37 min at zero cost. Both methods outperform our CLIP-based few-shot approach, which posts 0.580 accuracy, 0.560 sensitivity, 0.600 specificity, and an F1-score of 0.571 in 13.2 min.
Table 12: Performance and runtime comparison for Renault’s pipe staple. Bolded values denote the best-performing model for each metric.
Model Accuracy Sensitivity Specificity Precision F1 AUC Runtime (min)
CLIP Few-Shot 0.580 0.560 0.600 0.583 0.571 0.608 13.2
MobileNetV2-FS 0.750 0.720 0.780 0.766 0.742 0.590 1.4
AutoML 1 1 1 1 1 - - 98.0
- - indicates the value is not available/reported.
@@ 8.1 Dataset Description
Manual evaluation by human experts introduces significant limitations in industrial practice (20). The growing volume of material characterization data makes manual labeling a resource-intensive process. Expert analysis inherently contains subjective biases, which can limit the detection of subtle variations within natural data variance. Therefore, the materials science and manufacturing communities have increasingly adopted automated classification methods for microstructure analysis (11, 31). Our investigation of this case is motivated by the suitability of our proposed few-shot learning approach with CLIP to applications with limited training data. If successful, our proposed approach can automate the inspection of microstructure images to reduce the cost associated with expert-driven microstructural measurements.
We simulated EBSD microstructure images by combining random Voronoi tessellation with crystal orientation distribution functions (ODFs) (1). Voronoi edges represent grain size and morphology. The ODF assigns crystal orientations to grains based on predefined probability functions. Crystal orientations map to colors using standard EBSD measurement schemes. Figure 10 illustrates the four example simulated microstructures investigated in our study: (a) uniform (nominal) with random crystal orientations from a uniform distribution, (b) clustered orientation showing a band of grains with predominant crystal orientation surrounded by uniform ODF grains, (c) bimodal with crystal orientations from “bimodal” distribution, and (d) single crystal-like where crystals align toward one orientation. These classes reflect EBSD patterns in metal additive manufacturing. The uniform ODF represents nominal conditions. The other defective classes indicate defects resulting from variations in local or global heating and cooling dynamics during the manufacturing process.
@@
Our ODF simulations used weighted schemes to generate three severity levels per class (“low”, “medium”, “high”). These levels represent different degrees of deviation from the nominal uniform ODF. The simulated microstructural images are 400 x 400 pixels in size. For technical implementation details, we refer the interested reader to (37).
@@ 8.2 Few-Shot Classification Performance
Similar to case studies 1-4, Table 13 presents the performance of the few-shot learning for the binary image classification experiment (nominal vs. defective). Here, we define an image as defective if it belongs to any of the three primary defective classes (“band”, “bimodal”, and “single-crystal”). The few-shot learning binary classification results show that the model excelled at detecting defects with a sensitivity of 0.987, i.e., 222 out of the 225 defective images were correctly classified. Furthermore, nominal images were correctly classified as nominal 80.9% of the time, with 182 out of 225 nominal images classified correctly. Due to the balanced nature of the dataset, the overall accuracy represents the average of those two numbers and is approximately 90%.
@@
Given the relatively large number of potential defect classes (up to nine if we consider the class-severity combination), it would be interesting to investigate the model’s multiclass classification performance. Note that we can extract its multiclass classification performance without needing to “retrain” or “rerun” the model. Recall that our approach takes every test image and computes its cosine similarity with every image within the “nominal” and “defective” training dataset. From this comparison, we identify the most relevant image within each nominal and defective class, its match probability, and training-based captions (detailed image description) for its most similar training “nominal” and “defective” images. Consider our test set image microstructure_039.png to explain this concept further. The output from our CLIP-based implementation is shown in Figure 11(a). The Figure shows that the overall classification result is defective, as the probability of being defective (0.505) is slightly larger than that of being non-defective (0.495). We can also see that its most similar defective image is microstructure_001.png, which had a description/caption that ends with “low single crystal structure”. We extracted this description using simple text extraction strategies to obtain our multiclass results. We utilized two extraction strategies: (a) focusing only on whether it is “band”, “bimodal”, or “single-crystal”, and (b) assigning it to one of the nine defective sub-labels shown in Figure 10 by extracting severity (“low”, “medium”, and “high”) and the primary defect class (“band”, “bimodal”, and “single-crystal”) indicators.
@@
Figure 12: Confusion matrices and the corresponding classification performance metrics for primary and secondary multiclass labels for Case Study 5.
@@
The second row in Figure 12 captures the confusion matrix for the nominal class and the nine defective sublabels alongside their corresponding classification metrics table. As expected, with the increase in the classification degree of difficulty (the model must correctly identify both the primary defect class and its severity), the overall accuracy drops to 74.4%. However, compared to the primary classification example, the decrease in performance can only be attributed to misclassification errors among different severities of the same type of defect (crystal orientation) rather than errors between different crystal orientations.
@@
These results highlight the proposed approach’s capability to classify complex microstructure data through few-shot classification. The results also show the simplicity of extending our binary classification approach to multiclass classification without rerunning the model. This allows practitioners to perform classifications sequentially: first distinguishing between nominal and defective data and then identifying specific types of defects. Note that our results for this case study will likely be improved if we increase the number of learning examples, as shown in case studies 2 and 3. However, we do not explore this here for three reasons: (a) the obtained results demonstrate the approach’s potential and were satisfactory; (b) our primary rationale for presenting this case study was to show the ease of translating the binary classification results into a multiclass classification output, without the need to rerun or “retrain” the model; and (c) conciseness.
8.3 Comparison with Other Benchmark Approaches
A standard industrial practice for quantitative analysis and classification of microstructure images relies on computing simple synthetic descriptors. Typical descriptors relate to salient characteristics like grain size, morphology, and crystal orientation. We use two synthetic descriptors as a benchmark: (a) the entropy of RGB pixel intensities, and (b) the dominant color fraction. Their choice was motivated by the nature of simulated microstructures, which differ in terms of their Orientation Distribution Functions (ODFs). The RGB levels vary while the average grain and morphology distributions remain constant. These two descriptors capture microstructural heterogeneity (entropy) and the prevalence of any single crystallographic variant (dominant fraction) in two compact numbers. These are critical for texture quantification, grain statistics, and subsequent structure–property correlations (3, 39).
Each RGB microstructure was first converted to a grey-scale image. The distribution of grey levels was sampled with 256 equal-width bins to form a probability mass function, $p_{k}$. The Shannon entropy was then estimated as:
$H\;=\;-\sum_{k}p_{k}\,\log p_{k}$
For the dominant color fraction, all pixel colors were reshaped into an $N\times 3$ matrix in RGB space and partitioned into $K=5$ clusters using the $k-$means algorithm (with three random initializations, a maximum of 100 iterations). Being $N_{j}$ the number of pixels assigned to cluster $j$, the dominant-color fraction was defined as:
$f_{\mathrm{dom}}\;=\;\frac{\displaystyle\max_{1\leq j\leq K}N_{j}}{N}.$
The two descriptors were computed for every image. A k-NN classifier was trained and tested in the resulting bivariate space with $K=5$. The microstructure dataset was divided into training and test sets, analogously to the CLIP approach. This allowed us to evaluate performance against a benchmark representative of common industrial practice for microstructure image characterization and classification.
Table 14 captures the binary classification performance of our CLIP-based approach along with the three reference methods: (a) the common industrial practice of computing simple synthetic descriptors, (b) the MobileNetV2-based few-shot prototype, and (c) AutoML Vision. AutoML achieves the highest raw performance with 0.967 accuracy and 0.966. The descriptor-based benchmark trails substantially at 0.653 accuracy and 0.578 F1. Our few-shot CLIP approach and the MobileNetV2-based few-shot prototype outperform these industry-standard descriptors. They are competitive with the AutoML approach (especially since they incur zero costs and offer much shorter runtimes).
CLIP-based few-shot offers a unique advantage: its outputs support interpretable and hierarchical classification. Our implementation returns descriptive prompts for both nominal and defective reference images, as shown in Figure 11(a). These descriptions are generated without retraining the model. They enable practitioners to extend the binary classification setup to multiclass tasks using simple text parsing and matching techniques. The capacity for sequential classification without retraining or modifying the model underscores the flexibility of our few-shot approach for applications involving complex and nuanced visual structures.
Table 14: Performance and runtime comparison for microstructure images. Bolded values denote the best-performing model for each metric.
Model Accuracy Sensitivity Specificity Precision F1 AUC Runtime (min)
CLIP Few-Shot 0.898 0.987 0.809 0.838 0.906 0.952 34.4
Benchmark* 0.653 0.632 0.667 0.533 0.578 0.728 - -
MobileNet-FS 0.887 0.773 1.000 1.000 0.872 0.854 2.7
AutoML 0.967 0.991 0.942 0.991 0.966 - - 103.0
*Industrial Practice. - - indicates the value is not available/reported.
@@ 9.1 Insights from the Experiments
Our case studies demonstrate that CLIP, when adapted for few-shot learning, can serve as a powerful yet simple baseline for image-based quality control in a variety of contexts. Specifically, our CLIP applications demonstrate:
  • Strong Performance with Minimal Data: Most of our applications of the CLIP model achieved high accuracy with small learning sets–91% accuracy with 10 examples per class (metallic pan case), 97% accuracy with 50 examples per class (STS case), 90% accuracy with 72 examples (microarray structure case), and 80% accuracy with 350 examples per class in the low resolution ($85\times 85$ pixel) extrusion case. These are well below the $10,000+$ examples used in deep learning applications (e.g., see 18).
  • Competitive Performance with Implementation Advantages: CLIP demonstrated strong performance across diverse manufacturing scenarios with notable practical benefits. Performance varied by application complexity: it outperformed all methods in two case studies (metallic pan and STS), exceeded industry standards and slightly outperformed MobileNetV2-FS in microstructure classification, delivered competitive discriminative power (AUC 0.746 vs. 0.783) relative to MobileNetV2-FS in extrusion profile inspection (despite the below specification image size), but was significantly outperformed by AutoML in the complex pipe staple assembly task. Beyond predictive performance, CLIP offers substantial implementation advantages. It provides major advantages over AutoML in speed (9.1-34.4 minutes vs. 93.0-203.0 minutes), cost (zero vs. \$27.72 per case), user control, and interpretability. This combination positions CLIP as an attractive first baseline for practitioners seeking rapid deployment of image-based quality control systems.
@@
  • A Nonlinear Relationship between Learning Size and Performance: When increasing the learning sample size, some performance metrics improved rapidly for the first 50-100 examples per class with diminishing returns beyond 150-200 examples. The specific inflection point of the nonlinear improvement varied by application. In addition, performance gains across the size of the learning set were nonmonotonic. This phenomenon is likely due to the addition of learning set images, with some additional learning cases making it easier (or harder) to classify test images.
  • Robustness to Image Resolution: CLIP successfully processed images ranging from $250\times 250$ pixels (STS case) to full-resolution $3264\times 2448$ images (metallic pan case). While the ViT-L/14 model is designed for $336\times 336$ pixel inputs, it can effectively handle smaller and larger images through appropriate preprocessing. The impact of resolution on performance appears application-dependent, with optimal results generally achieved when image dimensions are larger or relatively close to the model’s design specifications. In the case of lower resolution images ($85\times 85$ pixels in the extrusion case), a larger training sample (350 images per class) was required to achieve adequate performance.
  • Sensitivity to Image Complexity: The model excelled with single-component images (metallic pan). Notably, CLIP performed well on STS images, a traditionally challenging domain for automated inspection. However, the CLIP model did not perform well with multi-component scenarios (pipe staple case). Our results suggest the transformer-based architectures may be well-suited for simple and texture-based tasks but less effective in scenarios requiring spatial reasoning across components. This limitation may stem from CLIP’s training data, which lacks industrial assembly images, and its default preprocessing, which can obscure peripheral or relational information critical to classification in complex assemblies.
@@ 9.3 Open Research Questions
9.4 Limitations
Our adaptation of the CLIP model demonstrates potential as a practical baseline for image inspection in manufacturing quality control. However, we acknowledge several limitations that should guide the future use of this method.
  • Preprocessing Constraints: CLIP requires square image inputs resized to fixed dimensions determined by the encoder model. Although cropping images is common practice in image analysis, it may exclude certain peripheral defects or distort spatial context.
  • No Use of Text Embeddings: Although CLIP supports image-text pairs in a shared embedding space, we did not leverage the text embeddings in our example. Our approach reflects real-world constraints in industrial vision systems where labeling images is costly. Thus, our approach may underutilize CLIP’s full capabilities.
  • Generalizability: Our five case studies span an array of quality control contexts; however, they do not represent the full diversity of industrial inspection tasks. By using the CLIP model as a baseline in future developments of application specific methodologies, we hope the understanding of the applicability of vision transformer models will expand.
@@ 10 Concluding Remarks
Our expository study demonstrates that our CLIP-based few-shot learning approach can serve as an effective baseline for image-based quality control in manufacturing. Through five case studies, we showed that CLIP can achieve high classification accuracy with relatively small learning sets, often requiring only 50-100 examples per class compared to the thousands typically needed for deep learning approaches. The microstructure case study further highlighted the model’s ability to adapt to multi-class classification tasks without retraining, demonstrating promising performance across multiple levels of classification granularity (from nominal vs. defective images to identifying specific defect types and their severities).
Our comprehensive comparison across all five case studies with MobileNetV2-based few-shot prototyping and AutoML Vision revealed that CLIP consistently delivered competitive performance while offering significant practical advantages in terms of implementation simplicity, runtime efficiency, and cost-effectiveness. Overall, the CLIP-based few-shot learning approach excelled at single-component inspection, texture analysis, and microstructure classification. However, its performance was poor in complex multi-component scenes, where more advanced computer vision and/or statistical monitoring approaches were needed. Our findings establish CLIP-based few-shot learning as a practical first baseline for quality engineers and researchers. It offers a simple yet powerful solution that should be evaluated before pursuing more complex implementations.
@@ Acknowledgement
The work of Hongyue Sun is partially supported by the NSF Grant No. FM-2134409 and CMMI-2412678. The research of Prof. Bianca Maria Colosimo and Prof. Marco Grasso was supported by ACCORDO Quadro ASI-POLIMI “Attività di Ricerca e Innovazione” n. 2018-5-HH.0, collaboration agreement between the Italian Space Agency and Politecnico di Milano.